Linda has some dimes nickels and quaters in a purse. The total value of these is $2.55. She has three times as many dimes as quarters . He has 3 more nickels than quarters. How many coin does she have of each type
Answer by stanbon!(97)

Let number of quarters be q.
Then number of dimes is 3q.
And number of nickels is q+3
Value of quarters is 25q cents
Value of dimes is 10(3q) = 30q cents
Value of nickels is 5(q+3) = 5x+15 cents
EQUATION
Value of nickels + Value of dimes + Value of quarters = 255 cents
5x+15 + 30q + 25q = 255
60q = 240
q = 4 (number of quarters)
3q = 12 (number of dimes)
q+3 = 7 (number of nickels)
Cheers,
Stan H.